Please help me with this problem:
The difference between an angle and 5/3 its complement is 18 degrees. Find the angles.

Answer by swincher4391(1107)

Let x be your angle.

When adding two angles together that are complementary, their sum is 90 degrees.
x - (5/3)(90-x) = 18
x - 150 + (5/3)x = 18
8/3 x = 168
x = 63
If one angle is 63, the other must be 27.